-6 + [-11 +5 ]/ 4^2 - (-9)=0
can anyone be kind enough to tell my why the answer to this equation equals 0

Answer by stanbon(75887)

-6 + [-11 +5 ]/ 4^2 - (-9)=0
---
-6 + [-6]/16 + 9 = 0
[-6/16] + 3 = 0
[-3/8] + 3 = 0
2 5/8 = 0
-----------------
It doesn't equal zero.
==========================
Your problem may be missing a variable, such as "x".
==========================
Cheers,
Stan H.